OpenStudy (anonymous):

the formula A= 1/2 bh represent the are of the triangle where a represnta the area, b is the base of the triangle and the h is the height of the triangle solve this formula for b.

OpenStudy (anonymous):

A=1/2bh 2(A=1/2bh)2 2A/h=bh/h b=2A/h

OpenStudy (anonymous):

thanks